OpenStudy (anonymous):

help me with the 4 questino pls

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2 can someone help me pls

OpenStudy (texaschic101):

45/16 = x/80 --- (45 goals to 16 games = x goals to 80 games) This is a proportion, so we must now cross multiply 16(x) = (45)(80) 16x = 3600 x = 3600/16 x = 225 The team would score 225 goals in 80 games questions ?

OpenStudy (anonymous):

d

OpenStudy (texaschic101):

yes...it is D